OpenStudy (anonymous):

simplify the expression using the rules for exponents (a^5b^-2c^-3)^4 / (a^-2b^-4c^5)^2 please help !

OpenStudy (whpalmer4):

\[\frac{(a^5b^{-2}c^{-3})^4}{(a^{-2}b^{-4}c^5)^2}\] I would start by using the property of exponents that \[(a^m)^n = a^{m*n}\] What do you get after doing that?

OpenStudy (anonymous):

I honestly have no idea , I don't really get this -.-

OpenStudy (whpalmer4):

go through the numerator. take each letter with its exponent, and use the rule I gave you. For starters: \((a^5)^4 = a^{5*4} = a^{20}\)

OpenStudy (whpalmer4):

next one up is \((b^{-2})^4 = \)

OpenStudy (a1234):

(a^5b^-2c^-3)^4 / (a^-2b^-4c^5)^2 (a^5(1/b^2)(1/c^3)^4 / ((1/a^2)(1/b^4)(c^5))^2 Can you simplify that further?

OpenStudy (whpalmer4):

Come on, go through the steps. \[(b^{-2})^4 = b^{-2*4} = \]What does \(-2*4=\)? That is the new exponent for \(b\). You're not allowed to say "idk" or "I don't get it" here.

OpenStudy (anonymous):

-8

OpenStudy (whpalmer4):

Very good. Now do the same for \[(c^{-3})^4 = \]

OpenStudy (anonymous):

okay so I understand you multiply the exponents , now I don't understand what to do after